# Rethinking Knowledge Distillation: A Data Dependent Regulariser With a Negative Asymmetric Payoff

### v1

@@ Appendix B Extended Functional Analysis: Information-Theoretic and Geometric Perspectives
| Activation Distance | 0.157 +- 0.001 | 0.157 +- 0.001 | 0.156 +- 0.001 | 0.155 +- 0.000 | 0.343 +- 0.000 | 0.581 +- 0.000 | 0.791 +- 0.000 |
| Rank Disagreement | 0.939 +- 0.000 | 0.939 +- 0.000 | 0.939 +- 0.000 | 0.939 +- 0.000 | 0.980 +- 0.000 | 0.984 +- 0.000 | 0.984 +- 0.000 |
| Prediction Disagreement | 0.153 +- 0.001 | 0.152 +- 0.001 | 0.151 +- 0.001 | 0.151 +- 0.001 | 0.190 +- 0.001 | 0.214 +- 0.000 | 0.324 +- 0.000 |
| JS Divergence | 0.040 +- 0.000 | 0.040 +- 0.000 | 0.039 +- 0.000 | 0.039 +- 0.000 | 0.171 +- 0.000 | 0.333 +- 0.000 | 0.533 +- 0.000 |
| Information Variation | 0.519 +- 0.017 | 0.520 +- 0.017 | 0.518 +- 0.022 | 0.533 +- 0.014 | 0.856 +- 0.002 | 0.897 +- 0.001 | 0.907 +- 0.002 |
| Procrustes Distance | 0.050 +- 0.000 | 0.050 +- 0.000 | 0.050 +- 0.000 | 0.049 +- 0.000 | 0.433 +- 0.000 | 0.664 +- 0.000 | 0.553 +- 0.000 |
| Accuracy | 0.605 +- 0.001 | 0.605 +- 0.000 | 0.604 +- 0.001 | 0.605 +- 0.001 | 0.607 +- 0.000 | 0.606 +- 0.001 | 0.580 +- 0.000 |
| Loss | 2.068 +- 0.001 | 2.065 +- 0.002 | 2.055 +- 0.001 | 2.043 +- 0.002 | 1.977 +- 0.001 | 2.497 +- 0.001 | 3.612 +- 0.002 |
Appendix C Feature Map Matching Knowledge Distillation
@@
The functional-similarity framework we introduce is agnostic to the form of teacher supervision: relation, feature, and contrastive approaches all deliver a teacher-derived signal that ultimately shapes the student’s output distribution. If a variant truly transfers richer or safer knowledge, it should manifest as higher functional similarity without the asymmetric amplification of teacher errors that we document.
To verify this, we run feature-map matching knowledge distillation (Romero et al., 2015) on the transformer model NanoGPT trained on Tiny Shakespeare. In this process, we try to align blocks in the transformers using Mean Squared Error (MSE) on the intermediate blocks’ outputs. We include this alignment in the backpropagation step111Feature-map matching knowledge distillation implementation: https://docs.pytorch.org/tutorials/beginner/knowledge_distillation_tutorial.html. We chose this dataset because it represents the case where standard knowledge distillation leads to the most significant negative asymmetric transfer.
When we run feature-map matching KD (Feature Map KD), we observe statistically significant knowledge transfer for blocks 4 and 5. Tables 11 and 12 report these results independently. However, we continue to observe asymmetric incorrect transfer, as shown in Figure 14(a). It is important to note that block 4 experiences less functional similarity transfer than block 5. As expected, this leads to less negative asymmetric transfer than observed for feature-map KD on block 5. The best accuracy is again recorded when using RCD for both blocks 4 and 5, but at a higher alpha value of 0.5, compared to the best results typically recorded for 0.1 with standard KD.

### v2

@@ 4.1.3 Language
| JS Divergence () | 0.053 | 0.049 | 0.037 | 0.032 | 0.067 | 0.192 | 0.449 |
| Accuracy () | 0.574 | 0.577 | 0.583 | 0.581 | 0.576 | 0.578 | 0.570 |
| Loss () | 1.559 | 1.542 | 1.496 | 1.500 | 1.507 | 1.839 | 2.995 |
For NanoGPT, KD provides statistically supported increases in functional similarity across all seeds, alongside improvements in accuracy and loss (Table 15). This shows that the RCD control does not yield the best performing model across all modalities. When considering the raw values for seed 0 (Table 14), KD yields a large deviation from the control baselines that scales with ; however, this is primarily driven by the largest error transfer that is recorded across all modalities (Figure 7). The maximum negative transfer is 10%, approximately larger than the positive transfer of just above 2%. Overall, these results confirm that across all major data modalities and architectures, negative asymmetric transfer is a prominent feature of KD: when knowledge transfer occurs, student capacity is preferentially used to match the teacher’s error patterns rather than its beneficial representations. In addition, we often observe that the RCD control can outperform KD, motivating stronger counterfactual baselines when interpreting KD gains.
| KD 0.1 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
| KD 0.5 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ |
| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✓✓ |
4.2 Standard Distillation: The Impact of Student Capacity on Negative Asymmetric Transfer
Our previous experiments in self-distillation provided an upper bound on knowledge transfer, revealing that negative asymmetric transfer often dominates. Now we study traditional knowledge distillation, where the student has less capacity than the teacher. Although this departs from our original setup, where the student can fully match the teacher, we use this setting to examine how transfer behaves between a larger teacher and a smaller student. A smaller student introduces uncertainty about whether limited capacity constrains knowledge transfer. However, knowledge distillation is typically used in this context, and to assess the generality of negative asymmetric transfer it is important to analyse the nature of knowledge transfer in this setting.
Our results for standard distillation show that our core insights from self-distillation extend to other distillation settings. These findings also reinforce the value of our initial setup for understanding knowledge transfer in KD. Apart from the student architecture’s implicit bias, which affects performance, there are no additional confounding factors influencing knowledge distillation in our experimental setup for standard distillation.
In the previous section, we demonstrated the generality of negative transfer across modalities in KD. Accordingly, here we focus on two common deep learning modalities: vision and language. We present results in Sections 4.2.1 and 4.2.2. These modalities exhibit distinct differences in both the proportion of knowledge transferred via KD and, consequently, the extent of negative asymmetric transfer. In vision, there is often limited knowledge transfer, apart from when teacher loss is high through the use of regularisation such as augmentation. In language, we observe statistically supported knowledge transfer alongside negative asymmetric transfer. In Section 4.2.1, we present results with ResNet50 teachers and ResNet18 students on Tiny ImageNet and ImageNet. In Section 4.2.2, we present results on Tinyshakespeare using Nano-GPT as the teacher and Pico-GPT as the student. Finally, in Section 4.2.3, we explore scaling laws for negative asymmetric transfer in KD as student capacity is gradually increased to have a granular insight on the emergence of knowledge transfer asymmetries.
4.2.1 Computer Vision
TinyImageNet
For this experiment, the ResNet50 teacher model was trained with stochastic gradient descent with a learning rate of 0.01 and a cosine annealing learning rate scheduler with T_max set to 100. It was trained for 100 epochs with a batch size of 256. The data was normalised with a mean of (0.485, 0.456, 0.406) and a standard deviation of (0.229, 0.224, 0.225). The ResNet18 student model was trained under the same conditions. Overall, we observe a low train loss for the teacher model (circa 0.0014) and a high train accuracy (0.9998); see Appendix Table 34. This low train loss corresponds to no statistically supported knowledge transfer across values; see Table 16, Appendix Table 35, 36, and Table 17. This result is consistent with the previous results and the intuition from the self-distillation experiment.
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
From Table 16, we observe that, compared to the SIDDO control, there is minimal functional similarity gain from using knowledge distillation across all similarity metrics. This indicates partial to no knowledge transfer, even when using a larger teacher. Furthermore, the best accuracy and loss are recorded for the RCD condition, highlighting that the regularisation effects of the KD objective can be more important for performance than knowledge transfer. Table 17 shows that knowledge transfer is largely not statistically supported across the majority of metrics when using the ResNet50 teacher with the ResNet18 student.
| KD 0.1 | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
| KD 0.5 | ✓✗✗ | ✗✗✗ | ✗✗✗ | ✓✓✗ | ✗✗✗ | ✗✗✗ |
| KD 0.9 | ✗✓✗ | ✗✗✗ | ✗✗✗ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
Analysing prediction agreement in Figure 8, we find that the student and teacher agree on fewer predictions than the baselines. Given the very limited functional transfer for this student–teacher pairing, this aligns with the expected low prediction agreement. Notably, although incorrect prediction agreement is not statistically supported, the student and teacher agree on incorrect predictions more than on correct predictions. This indicates that the small amount of knowledge transferred is still shaped by the negative asymmetric transfer highlighted earlier.
ImageNet
For this experiment, we use ImageNet with a ResNet50 teacher and a ResNet18 student (training details in Appendix C.2). For this dataset, we observe increases in functional similarity compared to the baselines in Table 18. However, only at at 0.9 do we obtain statistically supported transfer (Table 19) across the majority of functional similarity metrics. In line with our existing results, when there is statistically supported knowledge transfer from the teacher to the student, negative asymmetric transfer occurs with a bias towards teacher errors (Figure 9). Interestingly, while there is statistically supported transfer, for Rank disagreement the values recorded are very close in value and high indicating that for datasets with large classes rank order is more stochastic, which is somewhat expected. We record the largest relative deviations from the baselines for activation distance and prediction disagreement for this teacher seed. Largely, these results provide additional evidence that even on high complexity datasets such as ImageNet, negative asymmetric transfer remains prominent and increases with greater function transfer.
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.420 0.001 | 0.365 0.001 | 0.260 0.001 | 0.226 0.000 | 0.268 0.001 | 0.259 0.002 | 0.376 0.000 |
| Rank Disagreement | 0.997 0.000 | 0.997 0.000 | 0.997 0.000 | 0.997 0.000 | 0.997 0.000 | 0.997 0.000 | 0.997 0.000 |
| Prediction Disagreement | 0.264 0.003 | 0.256 0.002 | 0.239 0.002 | 0.235 0.002 | 0.259 0.001 | 0.274 0.002 | 0.308 0.002 |
| JS Divergence | 0.260 0.001 | 0.221 0.001 | 0.136 0.001 | 0.106 0.000 | 0.136 0.001 | 0.099 0.001 | 0.173 0.001 |
| Accuracy | 0.680 0.002 | 0.687 0.002 | 0.700 0.001 | 0.703 0.002 | 0.684 0.001 | 0.670 0.001 | 0.642 0.002 |
| Loss | 1.307 0.009 | 1.342 0.009 | 1.608 0.015 | 1.833 0.022 | 1.657 0.013 | 2.548 0.017 | 4.060 0.012 |
| KD 0.1 | ✗ | ✗ | ✗ | ✗ | ✓ | ✗ |
| KD 0.5 | ✗ | ✓ | ✓ | ✗ | ✓ | ✗ |
| KD 0.9 | ✓ | ✓ | ✓ | ✗ | ✓ | ✗ |
4.2.2 Language
The Nano-GPT teacher is a GPT2-style transformer configured with an embedding dimension of 384, a vocabulary size of 65, six attention heads, six transformer blocks, a dropout of 0.2, and a block size of 256. The Pico-GPT student differs by using a smaller embedding dimension of 192, halving the internal model width, while keeping all other hyperparameters identical to the teacher. The teacher and student are trained on TinyShakespeare, using the first 90% of the dataset for training and the last 10 % for testing. The dataset is tokenised with a character tokeniser, and models are trained autoregressively to predict the next character token. Training uses the Adam optimiser with a learning rate of , and a batch size of 64, for 5000 iterations. The student models share seeds and data orders from seeds 10 to 19 for the 10 runs used in averaging. This process is repeated for each of the three teachers trained with seeds 0 to 2.
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| KD 0.1 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✗ | ✗✗✗ |
| KD 0.5 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ |
| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ |
We observe a train loss for the teacher model of circa 0.86 with a high train accuracy (circa 0.72); see Appendix Table 38. We observe that this train loss corresponds to substantial knowledge transfer, which increases with ; see Tables 20 (Appendix Tables 39, 40) and Table 21. This substantial knowledge transfer coincides with an asymmetric payoff in prediction agreement that strongly favours incorrect predictions (Figure 10). These results match those in Section 4.1 and show that negative asymmetric transfer generalises to standard KD setups while showing that student capacity, as expected, limits the proportion of functional transfer.
4.2.3 Negative Transfer Scaling Laws
The previous sections established that when KD transfers knowledge, this transfer is negatively asymmetric. We now ask precisely how these effects evolve with capacity. Distillation Scaling Laws (DSL) (Busbridge et al., 2025) quantify how much student loss changes with compute, teacher quality, and model size. Our study complements DSL by asking how much is transferred as capacity grows: we decompose the distillation signal into correct vs. incorrect teacher–student agreement, offering a mechanistic interpretation of the “teacher quality” term and revealing negative-transfer regimes that are invisible from loss alone. Concretely, on TinyShakespeare we sweep student width from to in steps under a fixed-epoch budget matched to the teacher, using the same optimiser. For each width and , we measure the change in correct and incorrect agreement relative to the best control baseline (means SEM over 10 runs; teacher seed 0). Figure 11 shows three core trends:
1) Student capacity helps, but mainly by amplifying the teacher’s mistakes: as width increases, both correct and incorrect agreement rise, but incorrect agreement grows much faster (from to width at , correct agreement vs. incorrect ).
2) Small students suffer negative transfer: at - width, incorrect agreement is comparable to, or larger than, correct agreement.
3) Increasing capacity unlocks more of the distillation signal: however, what flows first, and most strongly, is the teacher’s error pattern.
Taken together, these scaling results calrify what drives the loss curves: KD acts as a data-dependent regulariser with a negative asymmetric payoff, and scaling up the student amplifies the asymmetry of knowledge transfer. As a result, we do not find that student capacity generally impacts the presence of negative asymmetric transfer in regimes where increased functional similarity occurs.
4.3 KD Variants: Feature-Map Matching Distillation
@@
The functional similarity framework we introduce is agnostic to the form of teacher supervision: relation-, feature-, and contrastive-based approaches all provide a teacher-derived signal that ultimately shapes the student’s output distribution. If a variant truly transfers richer or “safer” knowledge, it should manifest as higher functional similarity without the asymmetric amplification of teacher errors that we observe. To test this, we run feature-map matching KD (Romero et al., 2015) on the transformer model NanoGPT trained on TinyShakespeare. Specifically, we align transformer blocks using Mean Squared Error (MSE) on the intermediate block outputs, and include this alignment during backpropagation111Feature-map matching knowledge distillation implementation: https://docs.pytorch.org/tutorials/beginner/knowledge_distillation_tutorial.html. In this seciton, we focus on this dataset because it is the setting in which standard knowledge distillation exhibits the most pronounced negative asymmetric transfer.
@@
When we run feature-map matching KD, we observe statistically supported knowledge transfer for blocks 4 and 5, reported separately in Tables 22 and 23. Once again, we observe asymmetric incorrect transfer, as shown in Figure 12. Notably, block 4 exhibits less functional similarity transfer than block 5; as expected, this corresponds to less negative asymmetric transfer than observed for feature-map matching KD on block 5. The best accuracy is achieved using RCD for block 4 but at a higher value of 0.5, compared to the best results typically recorded at 0.1 for the RCD control, whereas for block 5 the best performance is recorded with KD at an of 0.5.